# Inexact Bregman Proximal Gradient Method and its Inertial Variant with Absolute and Partial Relative Stopping Criteria

### v4

@@ 1 Introduction
Compared to the PGM, the iterative scheme of the BPGM is more flexible. A suitable choice of the kernel function $\phi$ can help explore the underlying geometry or structure of the problem and potentially lead to a more tractable subproblem (1.2) (see, for example, [16, Section 5] and [56, Lemma 4]). More importantly, as investigated in recent insightful works [12, 50], the BPGM can be developed based on the notion of relative smoothness, which is weaker than the longstanding global Lipschitz gradient continuity assumption. This makes the BPGM applicable for a wider range of problems, such as the Poisson inverse problem [12], D-optimal design [50], symmetric nonnegative matrix factorization [33], orthogonal nonnegative matrix factorization [2], symmetric nonnegative matrix tri-factorization [1], deep linear neural networks [52], and sparse phase retrieval with Poisson loss [47], where $f$ involves a high-order polynomial function or a logarithm function in the form of relative entropy and log-determinant. Additionally, Ahookhosh and Nesterov [3] recently extended the results of [58] and shown that employing an (accelerated) inexact high-order proximal-point method at the upper level, while solving the proximal subproblem by the BPGM at the lower level, can result in a superfast method for solving problem (1.1). Due to this appealing potential, the BPGM and its variants have garnered increasing attention in recent years; see, for example, [4, 11, 12, 22, 34, 39, 50, 67, 70, 74].
@@
Unfortunately, although the underlying geometry/structure can possibly be captured by the kernel function, the subproblem (1.2) still has no closed-form solution in many applications and its computation may still be demanding numerically; see examples in [30, 43, 62] and the one in our numerical section. Thus, to make the BPGM implementable and practical, it must allow one to solve the subproblem approximately with progressive accuracy and the corresponding stopping criterion should be practically verifiable. However, unlike the inexact PGM which has been widely studied under different inexact criteria (see, for example, [32, 64, 71]), the study on the inexact BPGM is still at an early stage. Kabbadj [45] recently developed an inexact BPGM by allowing an $\varepsilon$-minimizer for each subproblem (1.2). Since the optimal objective value of subproblem (1.2) is generally unknown, such an inexact stopping criterion can be more difficult, if not impossible, to verify in practice. In [66], Stonyakin et al. proposed a general inexact algorithmic framework for first-order methods containing the BPGM as a special case, where an approximate solution of the subproblem (1.2) is accepted when a certain inexact condition [66, Definition 2.12] holds. In our context, their inexact condition can be covered by our absolute-type stopping criterion (AbSC) with the additional requirements that $\delta_{k}\equiv 0$ and $\bm{x}^{k+1}\equiv\widetilde{\bm{x}}^{k+1}$ (see Step 1 in Algorithm 1). However, this condition might be difficult to verify in some practical cases; see Remark 3.1 for more discussions. Moreover, both of the aforementioned two inexact conditions may not accommodate other types of errors that may occur when solving the subproblem. This limitation thus restricts their practical applications.
@@
Over the last few decades, Nesterov’s series of seminal works [53, 54, 55, 56] on accelerated gradient methods have inspired numerous accelerated variants of the PGM (see, e.g., [17, 23, 57, 68, 69]) as well as their inexact counterparts (see, e.g., [9, 18, 24, 44, 64, 71]). Naturally, those developments also motivated many researchers to explore whether and how the BPGM can be accelerated. A successful attempt was made by Auslender and Teboulle [10, Section 5], who proposed an improved interior gradient algorithm (a Bregman gradient scheme) for solving a special case of problem (1.1) with $P\equiv 0$ and showed that this algorithm can achieve a faster convergence rate of $\mathcal{O}(1/k^{2})$. Later on, it was extended by Tseng [68, 69] to handle problem (1.1) in a more general composite setting. Other accelerated variants of the BPGM can also be found in [46, 68, 69]. Note that all these works [10, 46, 68, 69] just mentioned require the assumptions that $f$ has a globally Lipschitz gradient and that the kernel function of the Bregman distance is strongly convex. Until recently, based on the relative smoothness assumption and a so-called triangle scaling property of the Bregman distance, Hanzely, Richtárik and Xiao [40] developed an accelerated BPGM that attains a convergence rate of $\mathcal{O}(1/k^{\gamma})$, where $\gamma\in(0,2]$ is the triangle scaling exponent. Later, a similar result was also derived under a slightly broader smoothness condition by Gutman and Peña [39, Section 3.3]. Very recently, Dragomir et al. [34] established a crucial result that the convergence rate of $\mathcal{O}(1/k)$ is indeed optimal for the BPGM to solve the class of problems merely satisfying the relative smoothness assumption, and thus the BPGM cannot be accelerated if no other regularity condition is imposed. But empirically improving the BPGM via an inertial step and a certain adaptive backtracking strategy is still possible, as discussed in [40, 51].
@@
However, among the aforementioned works that have investigated the possibility of improving the BPGM, none of them take into account of the possible errors incurred when one solves the involved subproblem approximately. Indeed, as far as we know, the rigorous study on the inexact inertial BPGM is very limited. Stonyakin et al. [66] have developed an inexact inertial variant of the BPGM in their work. However, as we have mentioned before, the inexact stopping criterion used in [66] can be covered by our absolute-type stopping criterion and may still require improvements regarding its practical verification. Moreover, we also note that Hien et al. [41] recently proposed a block alternating Bregman majorization-minimization framework with extrapolation, which gives an inertial BPGM if only one block of variables is considered, and employed a surrogate function for $P$ to obtain a more tractable auxiliary subproblem which possibly admits a closed-form solution. But such a surrogate function could be nontrivial to find when $P$ is complex, and the auxiliary subproblem needs to be solved exactly. Thus, the kind of inexactness considered in [41] is very different from what we consider in this paper. Moreover, replacing $P$ by its surrogate might also potentially slow down the BPGM, especially when the surrogate is not a good approximation of $P$.
@@
In this paper, to facilitate the practical implementations of the BPGM and its inertial variant, we attempt to develop their inexact counterparts and provide theoretical insights on how the error incurred in the inexact minimization of the subproblem would affect the convergence rate in terms of the objective function value. As we shall see later, our inexact framework is developed based on either an absolute-type stopping criterion or a partial relative-type stopping criterion, and both of them also have a two-point feature. Thus, the resulting inexact framework is rather broad and can handle different types of errors that may occur when solving the subproblem, and it is also helpful to circumvent the underlying feasibility difficulty in evaluating the Bregman distance when the problem has a complex feasible set; see Remark 3.1 for more details.
@@ 1.1 Contributions
    We develop an inexact Bregman proximal gradient method (iBPGM) based on either an absolute-type stopping criterion or a partial relative-type stopping criterion, both of which are distinct from the existing ones in [45, 66] and are practically verifiable. The $\mathcal{O}(1/k)$ convergence rate and the convergence of the sequence are also established for our iBPGM under some proper conditions.
@@
    We develop an inertial variant of our iBPGM (denoted by v-iBPGM) and establish the $\mathcal{O}(1/k^{\gamma})$ ($\gamma\geq 1$) convergence rate under an additional restricted relative smoothness assumption (see Assumption C). Our result can subsume the related results in [10, 40, 68, 69] wherein the subproblem is solved exactly, and as a byproduct, our analysis indeed provides a unified treatment of these existing results which are developed under different conditions. In particular, when the smooth part in the objective has a Lipschitz continuous gradient and the kernel function is strongly convex, we have $\gamma=2$ and thus the v-iBPGM improves the convergence rate of the iBPGM from $\mathcal{O}(1/k)$ to $\mathcal{O}(1/k^{2})$.
@@
    We conduct some preliminary numerical experiments to evaluate the performances of our iBPGM and v-iBPGM under different inexactness settings. The computational results empirically verify the necessity of developing inexact versions of those methods and also illustrate that different methods with different types of stopping criteria do have different inherent inexactness tolerance requirements, in accordance with our theoretical results.
@@ 1.2 Related works
where $\{\gamma_{k}\}$ is a given proximal parameter sequence, and $\{\eta_{k}\}$, $\{\nu_{k}\}$, $\{\mu_{k}\}$ are three given tolerance parameter sequences. By the definition of the Bregman distance, one can see that the subproblem (3.1) in the iBPGM can be equivalently reformulated as the form of (1.3) by setting $\psi:=\lambda\phi-f$ and $\gamma_{k}\equiv 1$. This demonstrates that, when the exact minimization of each subproblem is achievable, the BPGM can be viewed as a BPPA with a special kernel $\psi$ that depends on $f$. Such an equivalence has also been pointed out in [4, Theorem 4.1]. However, when considering the inexact minimization of each subproblem, this equivalence may no longer hold. Indeed, with $\psi:=\lambda\phi-f$, the inexact condition (1.4) used in the iBPPA cannot be equivalent to, nor implied by, condition (3.2) satisfying an absolute-type stopping criterion employed in our iBPGM. Consequently, the derivation and analysis in [31, 73] are not directly applicable to our iBPGM. In addition, we introduce a relative-type stopping criterion in this paper, which is not addressed in [31, 73]. As a result, by setting $f\equiv 0$, our iBPGM with this new relative-type stopping criterion also provides a new inexact version of the BPPA.
@@ 2 Notation and preliminaries
Assume that $f:\mathbb{E}\rightarrow(-\infty,\infty]$ is a proper closed convex function. For a given $\varepsilon\geq 0$, the $\varepsilon$-subdifferential of $f$ at $\bm{x}\in{\rm dom}\,f$ is defined by $\partial_{\varepsilon}f(\bm{x}):=\{\bm{d}\in\mathbb{E}:f(\bm{y})\geq f(\bm{x})+\langle\bm{d},\,\bm{y}-\bm{x}\rangle-\varepsilon,~{}\forall\,\bm{y}\in\mathbb{E}\}$, and when $\varepsilon=0$, $\partial_{\varepsilon}f$ is simply denoted by $\partial f$, which is referred to as the subdifferential of $f$. The conjugate function of $f$ is the function $f^{*}:\mathbb{E}\rightarrow(-\infty,\infty]$ defined by $f^{*}(\bm{y}):=\sup\left\{\langle\bm{y},\,\bm{x}\rangle-f(\bm{x}):\bm{x}\in\mathbb{E}\right\}$. A proper closed convex function $f$ is essentially smooth if (i) $\mathrm{int}\,\mathrm{dom}\,f$ is not empty; (ii) $f$ is differentiable on $\mathrm{int}\,\mathrm{dom}\,f$; (iii) $\|\nabla f(x_{k})\|\to\infty$ for every sequence $\{x_{k}\}$ in $\mathrm{int}\,\mathrm{dom}\,f$ converging to a boundary point of $\mathrm{int}\,\mathrm{dom}\,f$; see [61, page 251].
@@
Assumption A1 implies that the kernel function $\phi$ is of Legendre type; see [61, Page 258]. Moreover, $\overline{\mathrm{dom}}\,\phi=\mathcal{Q}$ implies $\mathrm{int}\,\mathrm{dom}\,\phi=\mathrm{int}\,\mathcal{Q}$ due to the convexity of $\mathrm{dom}\,\phi$ and [15, Proposition 3.36(iii)]. Then, one can see from Assumptions A2&3 that
@@ 3 An inexact Bregman proximal gradient method
In this section, we develop an inexact Bregman proximal gradient method (iBPGM) based on two types of inexact stopping criteria for solving problem (1.1) and study the convergence properties. The complete iterative framework is presented as Algorithm 1.
@@
    $\min\limits_{\bm{x}}\,P(\bm{x})+\langle\nabla f(\bm{x}^{k}),\,\bm{x}-\bm{x}^{k}\rangle+\lambda\mathcal{D}_{\phi}(\bm{x},\,\bm{x}^{k}),$ (3.1)
@@
One can see from Algorithm 1 that, at each iteration, our inexact framework allows one to approximately solve the subproblem (3.1) under condition (3.2) satisfying either an absolute-type stopping criterion (AbSC) or a partial [^1] Here, “partial” means that, unlike (AbSC), a non-zero error term $\Delta^{k}$ is not allowed in (ReSC). relative-type stopping criterion (ReSC). Since the subproblem (3.1) has an unique solution $\bm{x}^{k,*}\in\mathrm{dom}\,P\cap\mathrm{int}\,\mathrm{dom}\,\phi$ (by Assumption A5), then condition (3.2) always holds at $\bm{x}^{k+1}=\widetilde{\bm{x}}^{k+1}=\bm{x}^{k,*}$ and hence it is achievable. When setting $\varepsilon_{k}\equiv 0$ in (AbSC) or setting $\sigma=0$ in (ReSC), it means that $\bm{x}^{k+1}$ (equals to $\widetilde{\bm{x}}^{k+1}$) should be an optimal solution of the subproblem, and thus, our iBPGM readily reduces to the exact BPGM studied in [12, 50, 69, 74]. Moreover, when the conditions $\delta_{k}\equiv 0$ and $\bm{x}^{k+1}\equiv\widetilde{\bm{x}}^{k+1}$ are imposed in (AbSC), condition (3.2) with (AbSC) is equivalent to the inexact condition considered in [66, Definition 2.12] for the subproblem (3.1), and hence our iBPGM subsumes the inexact BPGM studied in [66] as a special case.
@@
Both inexact stopping criteria (AbSC) and (ReSC) are of two-point type (meaning that two points $\bm{x}^{k+1}$ and $\widetilde{\bm{x}}^{k+1}$ are used in the criterion), and they are inspired by the inexact stopping criterion proposed in our recent works [31, 73] for developing a new practical inexact Bregman proximal point algorithm. They may look unusual at the first glance, but actually provide a rather broad inexact framework in which the approximate solutions are handled through the error term $\Delta^{k}$ appearing on the left-hand-side of the optimality conditions, $\partial_{\delta_{k}}P$ (an approximation of $\partial P$), and the deviation $\mathcal{D}_{\phi}(\widetilde{\bm{x}}^{k+1},\,\bm{x}^{k+1})$. Thus, our inexact framework is adaptable to various types of errors that may incur when solving the subproblem inexactly. In particular, the admissible deviation allows $\partial_{\delta_{k}}P$ and $\nabla\phi$ to be computed at two slightly different points, respectively. Such a simple strategy would help to circumvent the potential feasibility difficulty of requiring $\bm{x}^{k+1}\in\mathrm{dom}\,P\cap\mathrm{dom}\,\nabla\phi$ by the exact BPGM studied in [12, 50, 69, 74] and by the inexact BPGM studied in [66], as exemplified and discussed in Remark 3.1. We have also noticed a recent work by Kabbadj [45], who proposed an inexact BPGM by allowing an $\varepsilon$-minimizer for the subproblem (3.1). However, such an inexact condition may be more difficult to verify in practice since the optimal objective value of the subproblem is generally unknown a priori. Moreover, it may not accommodate other types of errors that may occur when solving the subproblem. Finally, we would like to point out that both the pair $(\bm{x}^{k+1},\,\widetilde{\bm{x}}^{k+1})$ and the error pair $(\Delta^{k},\delta_{k})$ may not be unique. Any candidate satisfying condition (3.2) under either (AbSC) or (ReSC) is acceptable in our inexact framework.

### v5

@@ 1 Introduction
Compared to the PGM, the iterative scheme of the BPGM is more flexible. A suitable choice of the kernel function $\phi$ can help explore the underlying geometry or structure of the problem and potentially lead to a more tractable subproblem (1.2) (see, for example, [16, Section 5] and [56, Lemma 4]). More importantly, as investigated in recent insightful works [12, 50], the BPGM can be developed based on the notion of relative smoothness, which is weaker than the longstanding global Lipschitz gradient continuity assumption. This makes the BPGM applicable for a wider range of problems, such as the Poisson inverse problem [12], D-optimal design [50], symmetric nonnegative matrix factorization [33], orthogonal nonnegative matrix factorization [2], symmetric nonnegative matrix tri-factorization [1], deep linear neural networks [52], and sparse phase retrieval with Poisson loss [47]. Additionally, Ahookhosh and Nesterov [3] recently extended the results of [58] and shown that employing an (accelerated) inexact high-order proximal-point method at the upper level, while solving the proximal subproblem by the BPGM at the lower level, can result in a superfast method for solving problem (1.1). Due to this appealing potential, the BPGM and its variants have garnered increasing attention in recent years; see, for example, [4, 11, 12, 22, 34, 39, 50, 68, 71, 75].
@@
Unfortunately, although the underlying geometry/structure can possibly be captured by the kernel function, the subproblem (1.2) still has no closed-form solution in many applications and its computation may still be demanding numerically; see examples in [30, 43, 63] and the one in our numerical section. Thus, to make the BPGM truly implementable and practical, it must allow one to solve the subproblem approximately with progressive accuracy and the corresponding stopping criterion should be practically verifiable. However, unlike the inexact PGM which has been widely studied under different inexact criteria (see, for example, [32, 65, 72]), the study on the inexact BPGM is still at an early stage. Rebegoldi et al. [61] considered a more general problem of the form (1.1) by allowing a nonconvex $f$ and proposed a Bregman inexact linesearch–based forward–backward algorithm, where an $\varepsilon$-minimizer for the subproblem (1.2) is permitted at each iteration. However, they only establish the subsequential convergence for the proposed inexact algorithm. Moreover, their analysis requires the assumption that the kernel function $\phi$ satisfies the condition $0\in\mathrm{int}\,\mathrm{dom}\,\phi$. This assumption may exclude several commonly used kernel functions, such as the Boltzmann-Shannon entropy kernel function $\phi(\bm{x}):=\sum_{i=1}^{n}x_{i}(\log x_{i}-1)$ and the Burg’s entropy kernel function $\phi(\bm{x}):=-\sum_{i=1}^{n}\log x_{i}$. Subsequently, Kabbadj [45] developed an inexact BPGM that also permits an $\varepsilon$-minimizer for each subproblem (1.2) and established stronger convergence results under the relative smoothness assumption. However, since the optimal objective value of subproblem (1.2) is generally unknown, verifying an $\varepsilon$-minimizer may not be straightforward in practice. Recently, Stonyakin et al. [67] proposed a general inexact algorithmic framework for first-order methods containing the BPGM as a special case, where an approximate solution of the subproblem (1.2) is accepted when a certain inexact condition [67, Definition 2.12] holds. In our context, their inexact condition can be covered by our absolute-type stopping criterion (AbSC) with the additional requirements that $\delta_{k}\equiv 0$ and $\bm{x}^{k+1}\equiv\widetilde{\bm{x}}^{k+1}$ (see Step 1 in Algorithm 1). However, this condition might be difficult to verify in some practical cases; see Remark 3.1 for more discussions. In addition, all the aforementioned inexact algorithms in [45, 61, 67] do not accommodate the use of a relative-type stopping criterion, which limits their flexibility in practical applications.
@@
Over the last few decades, Nesterov’s works [53, 54, 55, 56] on accelerated gradient methods have inspired numerous accelerated variants of the PGM (see, e.g., [17, 23, 57, 69, 70]) as well as their inexact counterparts (see, e.g., [9, 18, 24, 44, 65, 72]). Naturally, those developments also motivated many researchers to explore whether and how the BPGM can be accelerated. A successful attempt was made by Auslender and Teboulle [10, Section 5], who proposed an improved interior gradient algorithm (a Bregman gradient scheme) for solving a special case of problem (1.1) with $P\equiv 0$ and showed that this algorithm can achieve a faster convergence rate of $\mathcal{O}(1/k^{2})$. Later on, it was extended by Tseng [69, 70] to handle problem (1.1) in a more general composite setting. Other accelerated variants of the BPGM can also be found in [46, 69, 70]. Note that all these works [10, 46, 69, 70] just mentioned require the assumptions that $f$ has a global Lipschitz gradient and that the kernel function of the Bregman distance is strongly convex. Until recently, based on the relative smoothness assumption and a so-called triangle scaling property of the Bregman distance, Hanzely, Richtárik and Xiao [40] developed an accelerated BPGM that attains a convergence rate of $\mathcal{O}(1/k^{\gamma})$, where $\gamma\in(0,2]$ is the triangle scaling exponent. Later, a similar result was also derived under a slightly broader smoothness condition by Gutman and Peña [39, Section 3.3]. Very recently, Dragomir et al. [34] established a crucial result that the convergence rate of $\mathcal{O}(1/k)$ is indeed optimal for the BPGM to solve the class of problems merely satisfying the relative smoothness assumption, and thus the BPGM cannot be accelerated if no other regularity condition is imposed. But empirically improving the BPGM via an inertial step and a certain adaptive backtracking strategy is still possible, as discussed in [40, 51].
@@
However, among the aforementioned accelerated works that have investigated the possibility of improving the BPGM, none of them take into account of the possible errors incurred when one solves the involved subproblem approximately. Indeed, as far as we know, the rigorous study on the inexact inertial BPGM is very limited. Stonyakin et al. [67] have developed an inexact inertial variant of the BPGM in their work. However, as we have mentioned before, the inexact stopping criterion used in [67] can be covered by our absolute-type stopping criterion and may still require improvements regarding its practical verification. Moreover, we also note that Hien et al. [41] recently proposed a block alternating Bregman majorization-minimization framework with extrapolation, which gives an inertial BPGM if only one block of variables is considered, and employed a surrogate function for $P$ to obtain a more tractable auxiliary subproblem which possibly admits a closed-form solution. But such a surrogate function could be nontrivial to find when $P$ is complex, and the auxiliary subproblem needs to be solved exactly. Thus, the kind of inexactness considered in [41] is very different from what we consider in this paper. Moreover, replacing $P$ by its surrogate might also potentially slow down the BPGM, especially when the surrogate is not a good approximation of $P$.
@@
In this paper, to facilitate the practical implementations of the BPGM and its inertial variant, we attempt to develop their inexact counterparts and provide theoretical insights on how the error incurred in the inexact minimization of the subproblem would affect the convergence rate in terms of the objective function value. As we shall see later, our inexact framework is developed based on either an absolute-type stopping criterion or a partial relative-type stopping criterion, and both of them also have a two-point feature. Thus, the resulting inexact framework is rather broad and can handle different types of errors that may occur when solving the subproblem, and it is also helpful to circumvent the underlying feasibility difficulty in evaluating the Bregman distance when dealing with the problem that has a complex feasible set; see Remark 3.1 for more details.
@@ 1.1 Contributions
    We develop an inexact Bregman proximal gradient method (iBPGM) based on either an absolute-type stopping criterion or a partial relative-type stopping criterion, both of which are distinct from the existing ones in [45, 61, 67] and are practically verifiable. The $\mathcal{O}(1/k)$ convergence rate and the convergence of the sequence are also established for our iBPGM under some proper conditions.
@@
    We develop an inertial variant of our iBPGM (denoted by v-iBPGM) and establish the $\mathcal{O}(1/k^{\gamma})$ ($\gamma\geq 1$) convergence rate under an additional restricted relative smoothness assumption (see Assumption C). Our result can subsume the related results in [10, 40, 69, 70] wherein the subproblem is solved exactly, and as a byproduct, our analysis indeed provides a unified treatment of these existing results which are developed under different conditions. In particular, when the smooth part in the objective has a Lipschitz continuous gradient and the kernel function is strongly convex, we have $\gamma=2$, and thus the v-iBPGM improves the convergence rate of the iBPGM from $\mathcal{O}(1/k)$ to $\mathcal{O}(1/k^{2})$.
@@
    We conduct preliminary numerical experiments to evaluate the performances of our iBPGM and v-iBPGM under different inexactness settings. The computational results empirically validate the necessity and significance of developing inexact versions for the BPGM and its inertial variant. Furthermore, they reveal that different methods with different types of stopping criteria have different inherent inexactness tolerance requirements, aligning with our theoretical findings.
@@ 1.2 Related works
where $\{\gamma_{k}\}$ is a given proximal parameter sequence, and $\{\eta_{k}\}$, $\{\nu_{k}\}$, $\{\mu_{k}\}$ are three given tolerance parameter sequences. By the definition of the Bregman distance, it can be easily verified that
$P(\bm{x})+f(\bm{x})+\mathcal{D}_{\psi}(\bm{x},\,\bm{x}^{k})=P(\bm{x})+f(\bm{x}^{k})+\langle\nabla f(\bm{x}^{k}),\,\bm{x}-\bm{x}^{k}\rangle+\lambda\mathcal{D}_{\phi}(\bm{x},\,\bm{x}^{k}),$
when $\psi:=\lambda\phi-f$ for any $\lambda>0$. Therefore, the subproblem (3.1) in the iBPGM developed later in Algorithm 1 is equivalent to the subproblem (1.3) with $\psi:=\lambda\phi-f$ and $\gamma_{k}\equiv 1$ in the above iBPPA. This demonstrates that, when the exact minimization of the subproblem (3.1) is achievable, the exact BPGM can be viewed as an exact BPPA with a special kernel $\psi$ that depends on $f$. Such an equivalence has also been pointed out in [4, Theorem 4.1]. However, when considering the inexact minimization of each subproblem, this equivalence may no longer hold. Indeed, with $\psi:=\lambda\phi-f$ and $\gamma_{k}\equiv 1$, the inexact condition (1.4) used in the above iBPPA becomes
$\Delta^{k}\in\partial_{\delta_{k}}(P+f)(\widetilde{\bm{x}}^{k+1})+\lambda\big{(}\nabla\phi(\bm{x}^{k+1})-\nabla\phi(\bm{x}^{k})\big{)}-\big{(}\nabla f(\bm{x}^{k+1})-\nabla f(\bm{x}^{k})\big{)}$
$~{}~{}\mathrm{with}~{}~{}\|\Delta^{k}\|\leq\eta_{k},~{}~{}\delta_{k}\leq\nu_{k},~{}~{}\mathcal{D}_{\lambda\phi-f}(\widetilde{\bm{x}}^{k+1},\,\bm{x}^{k+1})\leq\mu_{k},$
which is not equivalent to, nor implied by, condition (3.2) satisfying an absolute-type stopping criterion employed in our iBPGM, when $\widetilde{\bm{x}}^{k+1}\neq\bm{x}^{k+1}$. Consequently, the derivation and analysis in [31, 74] are not directly applicable to our iBPGM. In addition, we introduce a relative-type stopping criterion in this paper, which is not addressed in [31, 74]. As a result, by setting $f\equiv 0$, our iBPGM with this new relative-type stopping criterion also provides a new inexact version of the BPPA.
@@ 2 Notation and preliminaries
Assume that $f:\mathbb{E}\rightarrow(-\infty,\infty]$ is a proper closed convex function. For a given $\varepsilon\geq 0$, the $\varepsilon$-subdifferential of $f$ at $\bm{x}\in{\rm dom}\,f$ is defined by $\partial_{\varepsilon}f(\bm{x}):=\{\bm{d}\in\mathbb{E}:f(\bm{y})\geq f(\bm{x})+\langle\bm{d},\,\bm{y}-\bm{x}\rangle-\varepsilon,~{}\forall\,\bm{y}\in\mathbb{E}\}$, and when $\varepsilon=0$, $\partial_{\varepsilon}f$ is simply denoted by $\partial f$, which is referred to as the subdifferential of $f$. The conjugate function of $f$ is the function $f^{*}:\mathbb{E}\rightarrow(-\infty,\infty]$ defined by $f^{*}(\bm{y}):=\sup\left\{\langle\bm{y},\,\bm{x}\rangle-f(\bm{x}):\bm{x}\in\mathbb{E}\right\}$. A proper closed convex function $f$ is essentially smooth if (i) $\mathrm{int}\,\mathrm{dom}\,f$ is not empty; (ii) $f$ is differentiable on $\mathrm{int}\,\mathrm{dom}\,f$; (iii) $\|\nabla f(x_{k})\|\to\infty$ for every sequence $\{x_{k}\}$ in $\mathrm{int}\,\mathrm{dom}\,f$ converging to a boundary point of $\mathrm{int}\,\mathrm{dom}\,f$; see [62, page 251].
@@
Assumption A1 implies that the kernel function $\phi$ is of Legendre type; see [62, Page 258]. Moreover, $\overline{\mathrm{dom}}\,\phi=\mathcal{Q}$ implies $\mathrm{int}\,\mathrm{dom}\,\phi=\mathrm{int}\,\mathcal{Q}$ due to the convexity of $\mathrm{dom}\,\phi$ and [15, Proposition 3.36(iii)]. Then, one can see from Assumptions A2&3 that
@@ 3 An inexact Bregman proximal gradient method
In this section, we develop an inexact Bregman proximal gradient method (iBPGM) based on two types of inexact stopping criteria for solving problem (1.1) and study its convergence properties. The complete iterative framework is presented as Algorithm 1.
@@
    $\min\limits_{\bm{x}}~{}P(\bm{x})+\langle\nabla f(\bm{x}^{k}),\,\bm{x}-\bm{x}^{k}\rangle+\lambda\mathcal{D}_{\phi}(\bm{x},\,\bm{x}^{k}),$ (3.1)
@@
One can see from Algorithm 1 that, at each iteration, our inexact framework allows one to approximately solve the subproblem (3.1) under condition (3.2) satisfying either an absolute-type stopping criterion (AbSC) or a partial [^1] Here, “partial” means that, unlike (AbSC), a non-zero error term $\Delta^{k}$ is not allowed in (ReSC). relative-type stopping criterion (ReSC). Since the subproblem (3.1) has an unique solution $\bm{x}^{k,*}\in\mathrm{dom}\,P\cap\mathrm{int}\,\mathrm{dom}\,\phi$ (by Assumption A5), then condition (3.2) always holds at $\bm{x}^{k+1}=\widetilde{\bm{x}}^{k+1}=\bm{x}^{k,*}$ and hence it is achievable. When setting $\varepsilon_{k}\equiv 0$ in (AbSC) or setting $\sigma=0$ in (ReSC), it means that $\bm{x}^{k+1}$ (equals to $\widetilde{\bm{x}}^{k+1}$) should be an optimal solution of the subproblem, and thus, our iBPGM readily reduces to the exact BPGM studied in [12, 50, 70, 75]. Moreover, when the conditions $\delta_{k}\equiv 0$ and $\bm{x}^{k+1}\equiv\widetilde{\bm{x}}^{k+1}$ are imposed in (AbSC), condition (3.2) with (AbSC) is equivalent to the inexact condition considered in [67, Definition 2.12] for the subproblem (3.1), and hence our iBPGM subsumes the inexact BPGM studied in [67] as a special case.
@@
Both inexact stopping criteria (AbSC) and (ReSC) are of two-point type (meaning that two points $\bm{x}^{k+1}$ and $\widetilde{\bm{x}}^{k+1}$ are used in the criterion), and they are inspired by the inexact stopping criterion proposed in our recent works [31, 74] for developing a new practical inexact Bregman proximal point algorithm. They may look unusual at the first glance, but actually provide a rather broad inexact framework in which the approximate solutions are handled through the error term $\Delta^{k}$ appearing on the left-hand-side of the optimality conditions, $\partial_{\delta_{k}}P$ (an approximation of $\partial P$), and the deviation $\mathcal{D}_{\phi}(\widetilde{\bm{x}}^{k+1},\,\bm{x}^{k+1})$. Thus, our inexact framework is adaptable to various types of errors that may incur when solving the subproblem inexactly. In particular, the admissible deviation allows $\partial_{\delta_{k}}P$ and $\nabla\phi$ to be computed at two slightly different points, respectively. Such a simple strategy would help to circumvent the potential feasibility difficulty of requiring $\bm{x}^{k+1}\in\mathrm{dom}\,P\cap\mathrm{dom}\,\nabla\phi$ by the exact BPGM studied in [12, 50, 70, 75] and by the inexact BPGM studied in [67], as exemplified and discussed in Remark 3.1. We have also noticed the recent works by Rebegoldi et al. [61] and Kabbadj [45]. In their inexact Bregman-type algorithms, both approaches allow for an $\varepsilon$-minimizer for the subproblem (3.1) at each iteration. However, verifying an $\varepsilon$-minimizer may not be straightforward in practice, as the optimal objective value of the subproblem is generally unknown a priori. In addition, all these existing inexact algorithms in [45, 61, 67] do not accommodate the use of a relative-type stopping criterion, which limits their flexibility in practical applications. Finally, we would like to point out that both the pair $(\bm{x}^{k+1},\,\widetilde{\bm{x}}^{k+1})$ and the error pair $(\Delta^{k},\delta_{k})$ may not be unique. Any candidate satisfying condition (3.2) under either (AbSC) or (ReSC) is acceptable in our inexact framework.